# Socratic Reasoning Improves Positive Text Rewriting

### v1

@@ Abstract
Reframing a negative into a positive thought is at the crux of several cognitive approaches to mental health and psychotherapy that could be made more accessible by large language model-based solutions. Such reframing is typically non-trivial and requires multiple rationalization steps to uncover the underlying issue of a negative thought and transform it to be more positive. However, this rationalization process is currently neglected by both datasets and models which reframe thoughts in one step. In this work, we address this gap by augmenting open-source datasets for positive text rewriting with synthetically-generated Socratic rationales using a novel framework called SocraticReframe. SocraticReframe uses a sequence of question-answer pairs to rationalize the thought rewriting process. We show that such Socratic rationales significantly improve positive text rewriting for different open-source LLMs according to both automatic and human evaluations guided by criteria from psychotherapy research.
@@
Anmol Goel, Nico Daheim, Iryna Gurevych Ubiquitous Knowledge Processing Lab (UKP Lab) Department of Computer Science and Hessian Center for AI (hessian.AI) Technical University of Darmstadt www.ukp.tu-darmstadt.de
@@ 1 Introduction
Negative thoughts can have a profound effect on human judgement and well-being and often overshadow positive thoughts Vaish et al. (2008), because they can be emotionally deep-rooted and triggering when brought to surface Beck (1979). Cognitive Reframing is a highly-validated Cognitive Behavioral Therapy (CBT) intervention technique that aims to address this by identifying and reframing negative thoughts into positive ones Clark (2013) and has proven useful in both clinical therapeutic and self-help settings Williams (2001). For example, a negative thought “I am divorced. I will probably die soon and nobody will care for my children.” could be reframed to “I am divorced. I got out of an unhealthy relationship and can finally seek someone who makes me happier.” to provide a more positive outlook. Such techniques are not always accessible due to clinician shortages and a lack of self-help resources, but could be made more accessible using Large Language Models (LLMs).
Recent works in psychology have shown that the rationalization involved in breaking down negative thoughts into evidence and specific actions can be beneficial in overcoming negative feelings and thoughts McManus et al. (2012). This technique is also used in other therapeutic settings like Motivational Interviewing Miller and Rollnick (2012). One approach to rationalization is using guided discovery Overholser (2018). Guided Discovery is a collaborative method between a therapist and client that navigates important topics in multiple steps to improve rational thinking and adapt client behaviour. The Socratic method111The terminology for the Socratic method is not consistent within the literature Carey and Mullan (2004). In the context of this work, we use the term Socratic reasoning to refer to question-answer sequences that follow the Socratic method., which uses questioning strategies for uncovering a person’s beliefs and motivations, has shown to be a useful operationalization of this technique. Socratic reasoning typically uses probing questions to elicit alternative perspectives from the other person. Table 1 lists different types of Socratic questions that are common in literature and practice.
@@
Large Language Models (LLMs) hold promise to overcome the lack of clinician support and self-help resources by automated solutions. Notably, it has been shown that they can effectively guide text generation towards desired attributes like non-toxicity Liu et al. (2021); Zheng et al. (2023), style Reif et al. (2022); Tikhonov et al. (2019), or persona Song et al. (2020). While this makes them attractive for CBT-aware text rewriting with several public datasets being proposed recently (Ziems et al., 2022; Maddela et al., 2023; Sharma et al., 2023), they all fail to provide information on the rationalization process (see Figure 3), despite rationalization being pervasively used in CBT. This hinders model development and does not make the thought process behind reframing transparent.
Still, obtaining this information from humans can be costly. At the same time, step-by-step reasoning and rationalization with LLMs has progressed significantly outside of mental health applications. For example, models are used to generate thought chains to justify their reasoning process for solving complex tasks like math word problems (Kojima et al., 2022, inter alia). The scaffolded nature of Socratic questioning in guided discovery is analogous to such approaches to reasoning with LLMs. Therefore, we make use of rationales that are synthetically-generated from LLMs to augment CBT-aware text rewriting datasets with Socratic rationales. We show that this significantly improves positive text rewriting with LLMs that are finetuned using Socratic rationales.
Contribution
We propose SocraticReframe, a novel framework to improve LLM-based text-rewriting systems with synthetically-generated rationalization steps. We first enrich three existing datasets, namely CognitiveReframe Sharma et al. (2023), PatternReframe Maddela et al. (2023) and PositiveReframe Ziems et al. (2022) with such rationales. Then, we use them to finetune a series of language models for positive text rewriting. We find that models trained with Socratic rationales significantly outperform models trained without them. This is supported by both extensive automatic and human evaluation. Furthermore, we conduct a series of ablations which shows that the generated text is on average much more positive when rationales are used and validate their efficacy using information-theoretic measures.
Organization
The rest of the paper is organized as follows. Section 2 provides the background on Socratic questioning for cognitive reframing and NLP. We then introduce our framework, SocraticReframe in Section 3 followed by a discussion on the experimental setup and evaluation criteria in Section 4. The empirical results are discussed in Section 5 followed by human evaluations in Section 6. In Section 7, we conclude our study and discuss potential future directions based on our work, as well as limitations.
@@ 2 Background & Related Work
Clarification |
Questions to go deeper into a thought |
“Why do you say that?” |
Probing assumptions |
Questions to make someone think about unquestioned beliefs |
“What could we assume instead?” |
Probing reasons and evidence |
Questions digging into the reasoning behind a thought |
“How did you know that…?” |
Probing implications |
Questions probing the consequences of a thought |
“What is likely to happen if …?” |
Probing alternative viewpoints |
Questions about other, equally valid viewpoints |
“What is another way to look at it?” |
Question about the question |
Meta-questions about the question itself |
“Why do you think I asked this?” |
@@ 2.1 NLP and Mental Health
Research on NLP for mental health has primarily focused on utilizing linguistic features, as well as neural representations, to identify and analyze mental health conditions such as depression Rinaldi et al. (2020); Yates et al. (2017), anxiety Juhng et al. (2023); Wei et al. (2021), dyslexia Björnsdóttir et al. (2023); Gala and Ziegler (2016), autism Cho et al. (2022); Goodkind et al. (2018), and schizophrenia Mitchell et al. (2015); Sarioglu Kayi et al. (2017), among others. Typically, these studies rely on crowdsourced data or annotated social media posts to address the ethical and privacy concerns associated with medical data Harrigian et al. (2020); Moßburger et al. (2020); Turcan and McKeown (2019). Recently, efforts have been made to utilize datasets that are more representative of the interactions between a therapist and client in real world settings Pérez-Rosas et al. (2017); Shapira et al. (2022); Howes et al. (2014); Lee et al. (2019); Cao et al. (2019); Tanana et al. (2015); Shreevastava and Foltz (2021). Additionally, recent works have utilized synthetic data to enhance the performance of models in clinical Kazi and Kahanda (2019); Hiebel et al. (2023); Shim et al. (2021); Lindsay et al. (2022) and mental health Wu et al. (2023b) settings. Our work aims to contribute to the growing body of literature leveraging positive psychology Sheng et al. (2023) and LLMs to improve performance on mental health tasks.
@@ 2.2 Socratic Questioning
2.3 Reasoning with Large Language Models
Finally, related research has explored the abilities of LLMs to reason about complex problems. For example, Chain-of-Thought prompting (CoT) Wei et al. (2022) enables a model to generate intermediate reasoning steps before generating a final answer. Additionally, rationales, or reasoning paths, have been shown to improve the performance of LLMs when they are generated by either larger Shridhar et al. (2023) or even the same LLM Huang et al. (2023). Chen et al. (2023b) and Madaan et al. (2024) show evidence that LLMs can refine their own outputs to correct mistakes and generate the correct answer. Motivated by the prevalence of the Socratic method in psychology and the reasoning capabilities of LLMs, we leverage the synergy between the two and aim to improve LLM-based thought reframing by teaching them to verbalize a Socratic reasoning process.
@@ 3 SocraticReframe
Cognitive reframing aims to reframe a negative thought into a positive thought. We approach this with LLMs that generate a positive thought conditioned on the negative thought and metadata. However, directly training an LLM to map positive to negative thought does not make the process behind the reframing explicit, which is important both for interpretability and model performance. Our method SocraticReframe overcomes this by training the model to verbalize a Socratic rationale before generating the reframing. We introduce SocraticReframe in the following.
@@ 3.1 Generating Socratic Rationales
In particular, we few-shot prompt GPT-4 (Achiam et al., 2023) to generate Socratic question-answer pairs based on negative thought , positive thought , metadata , and in-context exemplars of Socratic rationales from clinical literature Padesky (1993) to ensure high quality and groundedness. The generated Socratic rationale then gives explicit reasoning steps for the particular reframing and can be used to augment the training data . This can be seen as a form of data augmentation and knowledge distillation into the model that is trained for cognitive reframing. As we will show in Section 6, the Socratic rationales are rated favourably by human annotators. An example rationale is shown in Figure 3. Details and examples are in Appendix A and Appendix D.
In addition to generating rationales, we prompt GPT-4 to classify the specific type of each generated Socratic question according to the six types described in Table 1, namely “Clarification”, “Probing Assumptions”, “Probing reasons or evidence”, “Probing implications”, “Probing alternative viewpoints”, and “Question”. These can be used for further insights into the reframing process.
@@ 4.1 Datasets
POSREF |
PATREF |
COGREF |
| Train |
6,679 |
5,249 |
400 |
| Test |
835 |
18,635 |
200 |
Persona |
Situation |
@@
| Content Preservation () | Transfer Strength () | ||||||||||||||||
| BLEU | BLEURT | Pos | Acc. (%) | ||||||||||||||
| Dataset | Model | FS | CoT | FT | SoC | FS | CoT | FT | SoC | FS | CoT | FT | SoC | FS | CoT | FT | SoC |
| LLaMa | 12.6 | 12.1 | 14.3 | 15.7 | 0.53 | 0.51 | 0.56 | 0.58 | 0.54 | 0.55 | 0.55 | 0.58 | 89.15 | 90.02 | 91.46 | 92.45 | |
| POSREF | Mistral | 13.1 | 13.3 | 15.1 | 15.9 | 0.55 | 0.55 | 0.58 | 0.61 | 0.49 | 0.48 | 0.51 | 0.52 | 90.10 | 91.16 | 91.10 | 92.09 |
| ChatGPT | 15.1 | 15.6 | - | - | 0.57 | 0.59 | - | - | 0.54 | 0.57 | - | - | 92.00 | 92.30 | - | - | |
| LLaMa | 71.9 | 72.0 | 72.3 | 75.2 | 0.64 | 0.64 | 0.66 | 0.69 | 0.81 | 0.80 | 0.83 | 0.85 | 97.02 | 97.23 | 97.67 | 97.90 | |
| PATREF | Mistral | 73.0 | 73.5 | 74.1 | 75.1 | 0.65 | 0.66 | 0.67 | 0.68 | 0.78 | 0.77 | 0.80 | 0.82 | 96.51 | 96.01 | 96.79 | 97.00 |
| ChatGPT | 73.8 | 73.9 | - | - | 0.63 | 0.65 | - | - | 0.81 | 0.83 | - | - | 97.61 | 97.60 | - | - | |
| LLaMa | 27.1 | 27.0 | 29.0 | 30.1 | 0.59 | 0.59 | 0.61 | 0.62 | 0.67 | 0.67 | 0.69 | 0.70 | 90.12 | 90.12 | 91.48 | 92.30 | |
| COGREF | Mistral | 27.9 | 27.5 | 30.0 | 31.7 | 0.61 | 0.60 | 0.62 | 0.63 | 0.68 | 0.69 | 0.70 | 0.72 | 90.94 | 91.67 | 91.88 | 92.56 |
| ChatGPT | 30.0 | 31.6 | - | - | 0.63 | 0.63 | - | - | 0.69 | 0.71 | - | - | 91.90 | 92.55 | - | - | |
@@ 4.2 Experimental Set-up
We test the models on different setups. The first two use prompting and the last two use finetuning:
Few Shot (FS): Few shot exemplars of negative and positive thought are given to the model which are handcrafted from literature on cognitive reframing.
Chain of Thought (CoT): The model is asked to think step-by-step and rationalize the reframed thought. 222We observed similar performance for zero-shot and few-shot CoT and therefore only report zero-shot CoT. We believe the similar performance in both setups stems from the nature of the reframing task where there is extra contextual information in the form of and the LLMs are able to generalize well even without few-shot exemplars.
Finetune (FT): The model is finetuned using LoRA to reframe a negative thought without a Socratic rationale.
Finetune with Socratic Rationales (SoC): The model is finetuned using LoRA to reframe a negative thought by first rationalising it using Socratic questioning and then generating the positive reframing. The prompt template used for finetuning is shown in Figure 8.

### v2

@@ Abstract
Reframing a negative into a positive thought is at the crux of several cognitive approaches to mental health and psychotherapy that could be made more accessible by large language model-based solutions. Such reframing is typically non-trivial and requires multiple rationalization steps to uncover the underlying issue of a negative thought and transform it to be more positive. However, this rationalization process is currently neglected by both datasets and models which reframe thoughts in one step. In this work, we address this gap by augmenting open-source datasets for positive text rewriting with synthetically-generated Socratic rationales using a novel framework called SocraticReframe. SocraticReframe uses a sequence of question-answer pairs to rationalize the thought rewriting process. We show that such Socratic rationales significantly improve positive text rewriting for different open-source LLMs according to both automatic and human evaluations guided by criteria from psychotherapy research. We validate our framework and the synthetic rationalizations with expert judgements from domain experts and psychology students in an IRB-approved annotation study. Our findings highlight the potential of utilizing the synergy between LLM reasoning and established psychotherapy techniques to build assistive solutions for reframing negative thoughts.
@@
Anmol Goela, Nico Daheima, Christian Montagb, Iryna Gurevycha aUbiquitous Knowledge Processing Lab (UKP Lab) Department of Computer Science and Hessian Center for AI (hessian.AI) Technical University of Darmstadt bCentre for Cognitive and Brain Sciences, Institute of Collaborative Innovation, University of Macau, Macau, China www.ukp.tu-darmstadt.de
@@ 1 Introduction
Negative thoughts can have a profound effect on human judgement and well-being. Oftentimes negative thoughts overshadow positive thoughts Vaish et al. (2008), because they can be emotionally deep-rooted and triggering when brought to surface Beck (1979). Cognitive Reframing is a highly-validated Cognitive Behavioral Therapy (CBT) intervention technique that aims to address this by identifying and reframing negative thoughts into positive ones Clark (2013) and has proven useful in both clinical therapeutic and self-help settings Williams (2001). Therefore, reframing holds great potential as a strategy for clinicians to provide help in overcoming negative thoughts. Beyond this, teaching people to reframe their own thoughts as a self-guided mental health intervention could provide them with a coping strategy that does not rely on clinical help (Jorm et al., 2006). However, coming up with effective reframes can be challenging and training people, both for clinical and self-help settings, to effectively reframe negative thoughts is a time-consuming and laborious process. Altogether, this poses a risk of mental health improvements being slowed by a lack of availability of knowledge and clinicians.
@@
Large Language Models (LLMs) hold great promise to overcome this lack of clinician support and self-help resources by providing assistive solutions and could also be used as a training resource. Notably, it has been shown that they can effectively guide text generation towards desired attributes like non-toxicity Liu et al. (2021); Zheng et al. (2023), style Reif et al. (2022); Tikhonov et al. (2019), or persona Song et al. (2020), and could therefore provide a diverse set of reframes. This motivates this work, where we aim to improve the reframing performance of LLMs.
While LLMs hold great promise for mental health applications due to their step-by-step reasoning and rationalization capabilities (Kojima et al., 2022, inter alia), they require careful investigation in tandem with domain experts. Several factors need to be taken into consideration before deploying LLM-based applications for client or therapist-facing applications. LLM-based solutions should not only be consistent with the client’s mental state, interpretable and avoid hallucinations, but importantly should also follow CBT guidelines.
Therefore, we aim to tackle this gap in literature by making the reasoning of LLMs explicit and include it as a supervision signal during training. We train LLMs for the positive text rewriting task by first generating a rationalization using the Socratic method111The terminology for the Socratic method is not consistent within the literature Carey and Mullan (2004). In the context of this work, we use the term Socratic reasoning to refer to question-answer sequences that follow the Socratic method., which uses questioning strategies for uncovering a person’s beliefs and motivations, and has shown to be a useful tool for CBT-based cognitive reframing. Socratic reasoning typically uses probing questions to elicit alternative perspectives. Table 1 lists different types of Socratic questions that are common in literature and practice. We call our method SocraticReframe.
We benchmark SocraticReframe with various state-of-the-art LLMs for the task of positive text rewriting using both automatic and human evaluations with clinical experts. Our results show that explicitly using Socratic reasoning as a training signal improves the text rewriting performance of LLMs while staying interpretable and faithful to CBT guidelines. We conduct extensive analysis on the synthetically-generated Socratic rationales and find that the socratic rationales are both informative to the model and adhere to clinical standards according to a known Socratic-questioning evaluation scheme. Here, the rationales are rated as highly helpful and relevant for the reframing process.
We use the generated rationales to augment three existing cognitive reframing datasets from related works. We release all data and code publicly to enable subsequent research in the mental health domain, for example, for clinician training.
@@ 2 Background & Related Work
Clarification | Questions to go deeper into a thought | “Why do you say that?” |
Probing assumptions | Questions to make someone think about unquestioned beliefs | “What could we assume instead?” |
Probing reasons and evidence | Questions digging into the reasoning behind a thought | “How did you know that…?” |
Probing implications | Questions probing the consequences of a thought | “What is likely to happen if …?” |
Probing alternative viewpoints | Questions about other, equally valid viewpoints | “What is another way to look at it?” |
Question about the question | Meta-questions about the question itself | “Why do you think I asked this?” |
@@ 2.1 NLP and Mental Health
Research on NLP for mental health has primarily focused on utilizing linguistic features, as well as neural representations, to identify and analyze mental health conditions such as depression Rinaldi et al. (2020); Yates et al. (2017), anxiety Juhng et al. (2023); Wei et al. (2021), dyslexia Björnsdóttir et al. (2023); Gala and Ziegler (2016), autism Cho et al. (2022); Goodkind et al. (2018), and schizophrenia Mitchell et al. (2015); Sarioglu Kayi et al. (2017), among others. Typically, these studies rely on crowdsourced data or annotated social media posts to address the ethical and privacy concerns associated with medical data Harrigian et al. (2020); Moßburger et al. (2020); Turcan and McKeown (2019). Recently, efforts have been made to utilize datasets that are more representative of the interactions between a therapist and client in real world settings Pérez-Rosas et al. (2017); Shapira et al. (2022); Howes et al. (2014); Lee et al. (2019); Cao et al. (2019); Tanana et al. (2015); Shreevastava and Foltz (2021). Additionally, recent works have utilized synthetic data to enhance the performance of models in clinical Kazi and Kahanda (2019); Hiebel et al. (2023); Shim et al. (2021); Lindsay et al. (2022) and mental health Wu et al. (2023b) settings. Chen et al. (2023b) propose Diagnosis-of-Thought prompting focusing only on diagnosing the type of cognitive distortion given a patient’s speech. In contrast, we develop a novel training framework for improving LLM performance on cognitive reframing. Our work aims to contribute to the growing body of literature leveraging positive psychology Sheng et al. (2023) and LLMs to improve performance on mental health tasks.
@@ 3 SocraticReframe
Cognitive reframing aims to reframe a negative thought into a positive thought. We approach this with LLMs that generate a positive thought conditioned on the negative thought and metadata. However, directly training an LLM to map positive to negative thought does not make the process behind the reframing explicit, which is important both for interpretability and model performance. Our method SocraticReframe overcomes this by training the model to verbalize a Socratic rationale before generating the reframing.
@@ 3.1 Generating Socratic Rationales
In particular, we few-shot prompt GPT-4 (Achiam et al., 2023) to generate Socratic question-answer pairs based on negative thought , positive thought , metadata , and in-context exemplars of Socratic rationales from clinical literature Padesky (1993) to ensure high quality and groundedness. The generated Socratic rationale then gives explicit reasoning steps for the particular reframing and can be used to augment the training data . This can be seen as a form of data augmentation and knowledge distillation into the model that is trained for cognitive reframing. As we will show in Section 6, the Socratic rationales are rated favourably by human annotators. An example rationale is shown in Figure 3. Details and examples are in Appendix A and Appendix D. In addition to generating rationales, we prompt GPT-4 to classify the specific type of each generated Socratic question according to the six types described in Table 1, namely “Clarification”, “Probing Assumptions”, “Probing reasons or evidence”, “Probing implications”, “Probing alternative viewpoints”, and “Question”. These can be used for further insights into the reframing process.
@@ 4.1 Datasets
POSREF | PATREF | COGREF | |
Train | 6,679 | 5,249 | 400 |
Test | 835 | 18,635 | 200 |
- | Persona | Situation |
@@
Content Preservation () | Transfer Strength () | |||||||||||||||
BLEU | BLEURT | Pos | Acc. (%) | Emp | ||||||||||||
Dataset | Model | LLaMa | Mistral | ChatGPT | LLaMa | Mistral | ChatGPT | LLaMa | Mistral | ChatGPT | LLaMa | Mistral | ChatGPT | LLaMa | Mistral | ChatGPT |
FS | 12.6 | 13.1 | 15.1 | 0.53 | 0.55 | 0.57 | 0.54 | 0.49 | 0.54 | 89.15 | 90.10 | 92.00 | 0.58 | 0.66 | 0.70 | |
CoT | 12.1 | 13.3 | 15.6 | 0.51 | 0.55 | 0.59 | 0.55 | 0.48 | 0.57 | 90.02 | 91.16 | 92.30 | 0.61 | 0.69 | 0.73 | |
POSREF | FT | 14.3 | 15.1 | - | 0.56 | 0.58 | - | 0.55 | 0.51 | - | 91.46 | 91.10 | - | 0.63 | 0.72 | - |
SoCCoT | 13.9 | 13.6 | 14.5 | 0.49 | 0.51 | 0.55 | 0.51 | 0.49 | 0.51 | 89.82 | 90.42 | 92.12 | 0.55 | 0.59 | 0.71 | |
SoC | 15.7 | 15.9 | - | 0.58 | 0.61 | - | 0.58 | 0.52 | - | 92.45 | 92.09 | - | 0.69 | 0.76 | - | |
FS | 71.9 | 73.0 | 73.8 | 0.64 | 0.65 | 0.63 | 0.49 | 0.78 | 0.81 | 97.02 | 96.51 | 97.61 | 0.81 | 0.80 | 0.87 | |
CoT | 72.0 | 73.5 | 73.9 | 0.64 | 0.65 | 0.65 | 0.48 | 0.77 | 0.83 | 97.23 | 96.01 | 97.60 | 0.83 | 0.81 | 0.89 | |
PATREF | FT | 72.3 | 74.1 | - | 0.66 | 0.67 | - | 0.51 | 0.80 | - | 97.67 | 96.79 | - | 0.89 | 0.85 | - |
SoCCoT | 70.2 | 73.5 | 71.2 | 0.59 | 0.61 | 0.62 | 0.48 | 0.62 | 0.78 | 95.02 | 95.64 | 96.10 | 0.80 | 0.82 | 0.86 | |
SoC | 75.2 | 75.1 | - | 0.69 | 0.68 | - | 0.52 | 0.82 | - | 97.90 | 97.00 | - | 0.90 | 0.88 | - | |
FS | 27.1 | 27.9 | 30.0 | 0.59 | 0.61 | 0.63 | 0.67 | 0.68 | 0.69 | 90.12 | 90.94 | 91.90 | 0.79 | 0.81 | 0.88 | |
CoT | 27.0 | 27.5 | 31.6 | 0.59 | 0.60 | 0.63 | 0.67 | 0.69 | 0.71 | 90.0 | 91.67 | 92.55 | 0.82 | 0.84 | 0.90 | |
COGREF | FT | 29.0 | 30.0 | - | 0.61 | 0.62 | - | 0.69 | 0.70 | - | 91.48 | 91.88 | - | 0.89 | 0.91 | - |
SoCCoT | 26.4 | 27.1 | 30.6 | 0.55 | 0.59 | 0.60 | 0.65 | 0.68 | 0.70 | 90.44 | 91.54 | 92.0 | 0.80 | 0.85 | 0.89 | |
SoC | 30.1 | 31.7 | - | 0.62 | 0.63 | - | 0.70 | 0.72 | - | 92.30 | 92.56 | - | 0.91 | 0.94 | - |
@@ 4.2 Experimental Set-up
We test the models on different setups. Few Shot (FS): Few shot exemplars of negative and positive thought are given to the model which are handcrafted from literature on cognitive reframing. Chain of Thought (CoT): The model is asked to think step-by-step and rationalize the reframed thought. 222We observed similar performance for zero-shot and few-shot CoT and therefore only report zero-shot CoT. We believe the similar performance in both setups stems from the nature of the reframing task where there is extra contextual information in the form of and the LLMs are able to generalize well even without few-shot exemplars. Finetune (FT): The model is finetuned using LoRA to reframe a negative thought without a Socratic rationale. Socratic CoT (SoCCoT): The typical CoT prompt is modified by replacing “think step-by-step” with an instruction to generate Socratic rationale before generating the reframe. Finetune with Socratic Rationales (SoC): The model is finetuned using LoRA to reframe a negative thought by first rationalising it using Socratic questioning and then generating the positive reframing. The prompt template used for finetuning is shown in Figure 9.